The Smith Trucking Company claims that the average weight of its delivery






trucks when fully loaded is 6000 pounds with a standard deviation of 120






pounds. 36 trucks are selected at random and their weights recorded.






Within what limits will the average weights of 90% of the 36 trucks lie?

1
Expert's answer
2022-04-24T18:04:55-0400

We have a normal distribution,

"\bar x=6000,\sigma=120, n=36."


The formula to calculate a confidence interval for a population mean is as follows:

"CI=\bar{x}\pm z\cdot\cfrac{\sigma}{\sqrt{n}},"

where:

  • "\bar{x}=6000-" sample mean
  • z- the chosen z-value, for a 90% confidence interval z = 1.645
  • "\sigma=120-" sample standard deviation
  • n =36 - sample size.


So,

"CI=6000 \pm 1.645\cdot\cfrac{120}{\sqrt{36}}=\\=6000\pm32.9=(5967.1, 6032.9)."